OpenStudy (anonymous):

don't know how to solve this -.- pls help me. And let me learn plss. log[10] 100^3

OpenStudy (***[isuru]***):

u want to simplify this ?

OpenStudy (anonymous):

yaa.. is the answer 8?

OpenStudy (***[isuru]***):

first... write 100 as a power of 10 ..what will u get ?

OpenStudy (anonymous):

\[\log_{10}100^3 \] \[=3\log_{10}100\] \[=3\log_{10}10^2\] \[=3*2\log_{10}10\] We know that \[\log_{10}10 = 1\] Therefore, \[=3*2\] \[=6\]

OpenStudy (anonymous):

@Japorized thanksss !!

OpenStudy (anonymous):

@bless: You're welcome :)